Williton News report up to 30th March 2013

The main occurrence during the last couple of weeks has been the departure of 08 shunter D3462 from the WSR.
Ian took what will probably have been the last picture of D3462 on WSR metals. Seen below, and taken on Sunday
24th March, it shows the loco positioned ready for dispatch to the Mid Hants Railway the following day.
D3462 is of course one of the ubiquitous 08 shunting locos that once numbered over a 1000 locos in one form or
other. They were extremely useful to BR, hence the large build numbers. Some are still useful to mainline operators
and many have found a busy retirement on heritage railways. The volunteers at Williton are somewhat saddened that
we have lost this useful loco that has given 16 years faithful service to the WSR while in our care. We would have
appreciated the opportunity to bid for it ourselves had we known it was available for sale.
D3462 is seen in the Loco Compound at Bishops Lydeard, waiting on the
departure road for the low-loader to arrive. Behind it, on the right, is new
arrival 09019, which belongs to 4160 Ltd and is currently not operational.
It had been hoped to do a main generator exam on D6566 on Saturday but it was just too cold to work inside the
loco for any length of time so this has been postponed until a warmer day! However, it was deemed OK to fit a
voltmeter into D6575 as this promised to be a much quicker job. It is installed adjacent to the Overspeed Device in
the Control Cubicle and is connected to control supply PA. As you can see in the picture below, the reading on
battery volts as shown is 103V and although this seemed a bit high, Ian confirmed this to be correct using a digital
voltmeter. Either this set of batteries are very good or they slipped in an extra cell that hasn't yet been noticed!
In comparison, Ian also measured the battery voltage on D6566 as 99.8V which is generally more typical and
reasonable after all the cold weather.
As part of the side project to build up a spare Triple Pump, Ian has re-assembled the previously seized fuel transfer
pump from the spare pump set that was acquired in 2011 and it now seems to work OK following stripping down
and cleaning.
Most importantly, the defective element from the cooker destined for D6575 has been sent away and await a quote
for repair/replacement.

The following week, Ian completed the re-assembly of the spare Triple Pump Motor.This was the original one fitted
to D6575 and had apparently been damaged in the fire so a replacement unit was fitted during the restoration of
the loco. After stripping down it was found that the problem was mainly smoke damage and so the inside of the
motor has been thoroughly cleaned and coated with anti-tracking varnish.
The brush gear will be re-fitted in due course following refurbishment and the motor can then be tested.
Roy is booked to perform the cleaning and painting of the motor exterior.
On Saturday 30th, a main generator examination was
performed on D6566. Ian led the proceedings and was
ably and variously assisted by Cameron, Dave, Brad and
Ed as a training exercise.
The connecting braids were disconnected and the brush
ring clamps slackened. Then the brushgear was rotated
around and the brush box insulators and busbars were
cleaned with electrical cleaning fluid. All brushes were
checked for length and free movement in the holders.

When the bottom brush box was revealed, it was found to
be covered in oil (as usual) with the result that the
brushes were seized in the holders, so it was removed
from the generator and taken outside the loco for
attention.
The brush box was cleaned and the brushes were freed.
Unfortunately two of the brushes were damaged during
the freeing operation and so had to be replaced with new
ones from stock. The brush box was then re-fitted to the
generator and the brush ring rotated back to the neutral
point and locked with the braids re-connected. That's
another job crossed off the list.
30th March was a big day for D9526. For the first time in many months, she was dragged out into the sunshine to
have her two bonnets refitted. Such is the recent progress of the rebuild that it was deemed right and proper to put
the lids back on the body to enable all their respective re-connections to take place.
In the Main Shed, recent work has concentrated on infrastructure maintenance. The 10 Ton crane has had its
annual service and some light repair work to get it through its insurance exam. While we have the use of a
scissor-lift platform, essential for work on the crane, the opportunity has been taken to access the lighting in the
roof and ensure it is safe and sound. The roof lights are now a mixture of mercury vapour and sodium lamps,
which give much better colour light balance. A defective light fitting was removed and replaced with an LED
floodlight as a trial. This has been deemed successful, and the decision has now been made to fit this type of
light wherever possible, which will help reduce our electricity bill!
More on page 2. Click here.